11 8 15 12 11 12 11 14 15 5 6 11

16 14 13 14 13 9 11 19 12 10 13 5

13 7 3 14 13 16 20 15 5 15 15 5

15 7 9 11 17 18 10 14 9 7 4 15

13 12 16 8 12 12


(a)  Find:

i.            Variance

ii. Standard deviation


1
Expert's answer
2021-08-03T13:15:22-0400

i. Variance

"var=\frac{\sum(xi-\bar{x})^2}{n-1}"

The table below gives "\sum(xi-\bar{x})^2"



From the table, "var=\frac{830}{53}=15.66"

ii standard deviation

"std=\sqrt{var}"

"=\sqrt{15.66}"

"=3.957"